OpenStudy (anonymous):

find the median to the nearest hundreth -1/3, 1.2, 0.48, -4/3, 0.75, 1.38

OpenStudy (anonymous):

Post the numbers from lowest to highest -4/3 , -1/3, .48, .75, 1.2, 1.38 Then find the middle number Since there are two add them together then divde by 2 .48 +.751.23 1.23 / 2 = .615

OpenStudy (anonymous):

sort them as \[-\frac{4}{3},-\frac{1}{3},0.48,0.75,1.2,1.38\] the 2 'middle' numbers would be \[0.48, 0.75\]so to find the median you would find the mean of those 2 numbers, which is \[\frac{0.48+0.75}{2}\]

OpenStudy (anonymous):

i only have 3 selections from my math problem they are 0.48 , -4/3 , 0.62 which outta those are the answer.

OpenStudy (anonymous):

.62 is the hundreth decimal place for .615